Factorise. 49x^2 –36 ???????????????????????????????????

 Factorise.
 49x^2 –36 ???????????????????????????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

49x^2–36 = (7a)^2-6^2
= (7a+6)(7a–6)
Using identity: x^2-y^2= (x+y)(x-y)

Thanks